# SO(3) real algebra method for finite baryon-number density QCD

### v1

@@ Abstract
For lattice QCD calculations at finite baryon-number densities, we propose the “SO(3) real algebra method”, in which the SU(3) gauge variable is divided into the SO(3) and SU(3)/SO(3) parts. In this method, we introduce the “maximal SO(3) gauge” by minimizing the SU(3)/SO(3) part of the SU(3) gauge variable. In the Monte Carlo calculation, the SO(3) real algebra method employs the SO(3) fermionic determinant, i.e., the fermionic determinant of the SO(3) part of the SU(3) gauge variable, in the maximal SO(3) gauge, as well as the positive SU(3) gauge action factor $e^{-S_{G}}$. Here, the SO(3) fermionic determinant is real, and it is non-negative for the even-number flavor case ($N_{f}=2n$) of the same quark mass, e.g., $m_{u}=m_{d}$. The SO(3) real algebra method alternates between the maximal SO(3) gauge fixing and Monte Carlo updates on the SO(3) determinant and $e^{-S_{G}}$. After the most importance sampling, the ratio of the SU(3) and SO(3) fermionic determinants is treated as a weight factor. If the phase factor of the ratio does not fluctuate significantly in the sampled gauge configurations, then lattice QCD calculations at finite densities would be feasible.
@@ I Introduction
Finite baryon-number density QCD is also important for understanding the QCD phase diagram [16, 17], nuclear systems, and neutron stars based on the fundamental principles of the strong interaction. However, a serious problem called the sign problem prevents practical calculations of SU(3) lattice QCD at finite densities [15]. In lattice QCD, many methods have been proposed to overcome the sign problem and to clarify the QCD phase diagram with a quark chemical potential $\mu$ and temperature $T$.
@@
For instance, the Taylor expansion method [18, 19] employs a Taylor expansion with respect to $\mu/T$ and is effective in regions of small chemical potentials, such as $\mu<T$. The complex Langevin method [20, 21, 22, 23] is one of the hopeful methods because its statistical sampling methods can handle complex interactions without encountering algorithmic difficulties and enable nonperturbative calculations, although the theoretical foundation for the convergence is incomplete and the correctness of solutions must be verified for each dynamical system [24, 25]. The reweighing method [26, 27] is a useful general approach to deal with controllable and difficult factors separately [28] and has been applied to finite density QCD [29]. An analytic continuation has been attempted from systems with pure imaginary chemical potentials $\mu\in i{\bf R}$ [30, 31, 32], where no sign problem appears. The Lefschetz thimble method [33] and its applications have been investigated as an interesting approach to high density QCD. This method uses a complex path where the imaginary part of the action is constant or optimized to control the sign problem [34, 35, 36, 37, 38]. For heavy quark systems, a sign-problem-free effective model constructed by combining strong coupling and hopping parameter expansion has been investigated as an indirect approach to finite density QCD [39]. Despite the development of calculation techniques mentioned above, no conclusive method has yet been established.
@@
Instead, many authors have studied systems that resemble finite density QCD without the sign problem. For instance, since SU(2) is pseudo-real, SU(2) color QCD does not encounter the sign problem, and high-density systems have been studied in lattice QCD [41, 40, 42, 43, 44], although baryons are bosons in SU(2) QCD. Even in SU(3) QCD, no sign problem appears for the systems with isospin chemical potentials [45, 46] for the $N_{f}=2$ case with the same u, d quark mass, and lattice QCD simulations have been performed [47]. This type of system could be found in the mismatch of the proton-to-neutron ratio, as is the case inside neutron stars. Also, in SU(3) QCD, no sign problem appears for systems with a chiral chemical potential [48] as a result of the chiral magnetic effect [49]. Such a system may be realized in the presence of strong magnetic fields [50, 51, 52]. Although these studies may provide insight into the qualitative behavior of high density QCD, these examples are rather special or differ significantly from real finite-density SU(3) color QCD.
@@ II Finite baryon-number density QCD
In finite density SU(3) QCD, although the generating functional $Z$ itself is real, the phase factor of the fermionic determinant fluctuates violently and cancels itself out, making numerical simulations impossible. This is the sign problem.
However, if the phase factor is sufficiently suppressed during the most importance sampling, Monte Carlo simulations could be performed. Note again that the generating functional $Z$ itself is real, and therefore the sampled gauge configurations can, in principle, be chosen so that their determinant is close to being real.
@@ VII SO(3) real algebra method in SU(3) QCD
Our strategy is to divide the SO(3) part, which is essentially real, from the SU(3) fermionic determinant in the maximal SO(3) gauge. The QCD generating functional in the maximal SO(3) gauge at finite densities is expressed as
$Z_{\rm SO(3)}(\mu)$ $=$ $\int DAe^{-S_{G}[A]}~\delta^{\rm GF}_{\rm SO(3)}[A]~\Delta_{\rm SO(3)}^{\rm FP}[A]~{\rm Det}(\not\!\!D+m+\mu\gamma_{4})$ (146)
@@
$=$ $\sum_{n=1}^{\infty}\frac{(-)^{n+1}g^{2n}}{2n}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n}+i\sum_{n=1}^{\infty}\frac{(-)^{n}g^{2n+1}}{2n+1}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n+1}.$ (158)
@@
${\rm ln~}\frac{{\rm Det}(\not\!\!D+m+\mu\gamma_{4})}{{\rm Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})}$ $=$ $p+iq,$ (165)
${\rm i.e.,}\quad\frac{{\rm Det}(\not\!\!D+m+\mu\gamma_{4})}{{\rm Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})}$ $=$ $e^{p+iq},$ (166)
where $p$ and $q$ are the real numbers defined by
$p\equiv\sum_{n=1}^{\infty}\frac{(-)^{n+1}g^{2n}}{2n}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n}\in{\bf R},$ (167)
$q\equiv\sum_{n=1}^{\infty}\frac{(-)^{n}g^{2n+1}}{2n+1}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n+1}\in{\bf R}.$ (168)
For the fermionic determinant ratio, the real number $p$ yields a harmless real factor of $e^{p}$, and $q$ gives the phase factor $e^{iq}$, which has the potential to cause the sign problem in the Monte Carlo method.
If the norm of $A^{\mu}_{\rm SU(3)/SO(3)}$ is well suppressed in the maximally SO(3) gauge, $q$ may not fluctuate significantly over the sampled gauge configurations, and also the infinite series in Eq. (LABEL:eq:expansion) could be truncated to a finite one. Even after the truncation, the effect of the chemical potential $\mu$ is firmly included in the SO(3) fermionic determinant.
@@ VIII SO(3) real algebra method
Now, we formulate the SO(3) real algebra method in SU(3) lattice QCD at a finite density with a quark-number chemical potential $\mu$. Using the SU(3) lattice fermion kernel $K[U;\mu]$, the SU(3) fermionic determinant is expressed as ${\rm Det}~K[U;\mu]$. The lattice QCD generating functional in the maximal SO(3) gauge at finite densities is expressed as
$Z_{\rm SO(3)}(\mu)$ $=$ $\int DUe^{-S_{G}[U]}~\delta^{\rm GF}_{\rm SO(3)}[U]~\Delta_{\rm SO(3)}^{\rm FP}[U]~{\rm Det}~K[U;\mu]$ (169)
@@
The next important step is to check the phase factor of the determinant ratio ${\rm Det}~K[U_{k};\mu]/{\rm Det}~K[u_{k};\mu]$ in the maximal SO(3) gauge in an actual lattice QCD Monte Carlo simulation. The degree of fluctuation would depend on the lattice volume, the chemical potential $\mu$, the fermion mass $m$ and so on. If there is a window that is close to the physical situation, this approach would be realistic for exploring finite density QCD.
@@ IX Summary and conclusion
In this paper, we have proposed the SO(3) real algebra method for lattice QCD calculations at finite baryon-number densities. In this method, we have divided the SU(3) gauge variable into the SO(3) and SU(3)/SO(3) parts. Then, we have introduced the maximal SO(3) gauge by minimizing the SU(3)/SO(3) part of the SU(3) gauge variable. We have considered the SO(3) fermionic determinant, that is, the fermionic determinant of the SO(3) part of the SU(3) gauge variable.
@@
The SO(3) real algebra method alternates between the maximal SO(3) gauge fixing and Monte Carlo updates on the SO(3) determinant and $e^{-S_{G}}$. After the most importance sampling, we have treated the remaining ratio of the SU(3) and SO(3) fermionic determinants as a weight factor in the reweighting method. In fact, we have only used gauge fixing for the most importance sampling, and the final result is fully gauge invariant. If the phase factor of the ratio does not fluctuate significantly in the sampled gauge configurations, then lattice QCD calculations at finite densities would be feasible. This method can be extended to finite temperature QCD by imposing temporal (anti)periodicity. In fact, this method can be used to study QCD at finite densities and temperatures.
@@
If there is a window that is close to the physical situation, this approach would be realistic for lattice calculations of QCD at finite densities and temperatures, and the QCD phase diagram would be obtained.
@@
Compared to other methods, the SO(3) real algebra method is a type of reweighting method [26, 27, 28]. This method is philosophically similar to the Lefschetz thimble method [33, 34, 35, 36, 37, 38], which reduces phase fluctuations by taking a suitable complex path, while this method reduces them by taking a suitable gauge. Unlike the Taylor expansion method [18, 19], this method does not use an expansion by $\mu$, so its applicable region is not limited to small $\mu$. Even if this method alone may not sufficiently suppress phase fluctuations, combining it with the other methods mentioned in Sec. I may enable the numerical simulation of finite density QCD feasible.

### v2

# SO(3) real algebra method for SU(3) QCD at finite baryon-number densities

@@ Abstract
For SU(3) lattice QCD calculations at finite baryon-number densities, we propose the “SO(3) real algebra method”, in which the SU(3) gauge variable is divided into the SO(3) and SU(3)/SO(3) parts. In this method, we introduce the “maximal SO(3) gauge” by minimizing the SU(3)/SO(3) part of the SU(3) gauge variable. In the Monte Carlo calculation, the SO(3) real algebra method employs the SO(3) fermionic determinant, i.e., the fermionic determinant of the SO(3) part of the SU(3) gauge variable, in the maximal SO(3) gauge, as well as the positive SU(3) gauge action factor $e^{-S_{G}}$. Here, the SO(3) fermionic determinant is real, and it is non-negative for the even-number flavor case ($N_{f}=2n$) of the same quark mass, e.g., $m_{u}=m_{d}$. The SO(3) real algebra method alternates between the maximal SO(3) gauge fixing and Monte Carlo updates on the SO(3) determinant and $e^{-S_{G}}$. After the most importance sampling, the ratio of the SU(3) and SO(3) fermionic determinants is treated as a weight factor. If the phase factor of the ratio does not fluctuate significantly among the sampled gauge configurations for a set of parameters (e.g., volume, chemical potential, and quark mass), then SU(3) lattice QCD calculations at finite densities would be feasible.
@@ I Introduction
Finite baryon-number density QCD is also important for understanding the QCD phase diagram [16, 17, 18, 19], nuclear systems, and neutron stars based on the fundamental principles of the strong interaction. However, a serious problem called the sign problem prevents practical calculations of SU(3) lattice QCD at finite densities [15]. In lattice QCD, many methods have been proposed to overcome the sign problem and to clarify the QCD phase diagram with a quark chemical potential $\mu$ and temperature $T$.
@@
For instance, the Taylor expansion method [20, 21] employs a Taylor expansion with respect to $\mu/T$ and is effective in regions of small chemical potentials, such as $\mu<T$. The complex Langevin method [22, 23, 24, 25] is one of the hopeful methods because its statistical sampling methods can handle complex interactions without encountering algorithmic difficulties and enable nonperturbative calculations, although the theoretical foundation for the convergence is incomplete and the correctness of solutions must be verified for each dynamical system [26, 27]. The reweighing method [28, 29] is a useful general approach to deal with controllable and difficult factors separately [30] and has been applied to finite density QCD [31]. An analytic continuation has been attempted from systems with pure imaginary chemical potentials $\mu\in i{\bf R}$ [32, 33, 34, 35], where no sign problem appears. The Lefschetz thimble method [36] and its applications have been investigated as an interesting approach to high density QCD. This method uses a complex path where the imaginary part of the action is constant or optimized to control the sign problem [37, 38, 39, 40, 41]. For heavy quark systems, a sign-problem-free effective model constructed by combining strong coupling and hopping parameter expansion has been investigated as an indirect approach to finite density QCD [42]. The complex phase of the fermionic determinant in finite density QCD has been analyzed in the large $N_{c}$ context [43]. Despite the development of calculation techniques mentioned above, no conclusive method has yet been established.
@@
Instead, many authors have studied systems that resemble finite density QCD without the sign problem. For instance, since SU(2) is pseudo-real, SU(2) color QCD does not encounter the sign problem, and high-density systems have been studied in lattice QCD [45, 44, 46, 47, 48], although baryons are bosons in SU(2) QCD. Even in SU(3) QCD, no sign problem appears for the systems with isospin chemical potentials [32, 49, 50] for the $N_{f}=2$ case with the same u, d quark mass, and lattice QCD simulations have been performed [51]. This type of system could be found in the mismatch of the proton-to-neutron ratio, as is the case inside neutron stars. Also, in SU(3) QCD, no sign problem appears for systems with a chiral chemical potential [52] as a result of the chiral magnetic effect [53]. Such a system may be realized in the presence of strong magnetic fields [54, 55, 56]. Although these studies may provide insight into the qualitative behavior of high density QCD, these examples are rather special or differ significantly from real finite-density SU(3) color QCD.
@@ II Finite baryon-number density QCD
Just as the partition function is real in a usual statistical system, the generating functional $Z(\mu)$ is also real. In finite density SU(3) QCD, although $Z(\mu)$ itself is real, the fermionic determinant is generally complex, and its phase factor fluctuates violently and cancels out, making numerical simulations impossible. This is the sign problem. However, if the phase factor is sufficiently suppressed during the most importance sampling, Monte Carlo simulations could be performed. Note again that the generating functional $Z(\mu)$ itself is real, and therefore the sampled gauge configurations can, in principle, be chosen so that their determinant is close to being real.
@@ VII SO(3) real algebra method in SU(3) QCD
Our strategy is to divide the SO(3) part, which is essentially real, from the SU(3) fermionic determinant in the maximal SO(3) gauge. The SU(3) QCD generating functional in the maximal SO(3) gauge at finite densities is expressed as
$Z^{\rm~SU(3)}_{\rm SO(3)~gauge}(\mu)$ $=$ $\int DAe^{-S_{G}[A]}~\delta^{\rm GF}_{\rm SO(3)}[A]~\Delta_{\rm SO(3)}^{\rm FP}[A]~{\rm Det}(\not\!\!D+m+\mu\gamma_{4})$ (146)
@@
$=$ $\sum_{n=1}^{\infty}\frac{(-)^{n+1}g^{2n}}{2n}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n}+i\sum_{n=0}^{\infty}\frac{(-)^{n}g^{2n+1}}{2n+1}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n+1}.$ (158)
@@
${\rm ln~}\frac{{\rm Det}(\not\!\!D+m+\mu\gamma_{4})}{{\rm Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})}$ $=$ $p(\mu)+iq(\mu),$ (165)
${\rm i.e.,}\quad\frac{{\rm Det}(\not\!\!D+m+\mu\gamma_{4})}{{\rm Det}(\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4})}$ $=$ $e^{p(\mu)+iq(\mu)},$ (166)
where $p(\mu)$ and $q(\mu)$ are the real numbers defined by
$p(\mu)\equiv\sum_{n=1}^{\infty}\frac{(-)^{n+1}g^{2n}}{2n}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n}\in{\bf R},$ (167)
$q(\mu)\equiv\sum_{n=0}^{\infty}\frac{(-)^{n}g^{2n+1}}{2n+1}{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{2n+1}\in{\bf R}.$ (168)
For the fermionic determinant ratio, the real number $p(\mu)$ yields a harmless real factor of $e^{p(\mu)}$, and $q(\mu)$ gives the phase factor $e^{iq(\mu)}$, which has the potential to cause the sign problem in the Monte Carlo method.
Using the real nature of Eq. (164) and the anti-hermitian nature of $\not\!\!D_{\rm SO(3)}$, we find
${\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m-\mu\gamma_{4}}\right)^{n}$ (169)
$=$ ${\rm Tr}\left[\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m-\mu\gamma_{4}}\right)^{n}\right]^{\dagger}$ (170)
$=$ ${\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{-\not\!\!D_{\rm SO(3)}+m-\mu\gamma_{4}}\right)^{n}$ (171)
$=$ ${\rm Tr}\left[\gamma_{5}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{-\not\!\!D_{\rm SO(3)}+m-\mu\gamma_{4}}\right)^{n}\gamma_{5}\right]$ (172)
$=$ $(-)^{n}~{\rm Tr}\left(\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}\right)^{n}.$ (173)
Then, $p(\mu)$ and $q(\mu)$ are even and odd functions of $\mu$, respectively,
$p(-\mu)=p(\mu),\quad q(-\mu)=-q(\mu).$ (174)
This leads to
$q(\mu=0)=0,$ (175)
meaning that the fermionic determinant ratio is real at zero density, $\mu=0$. Using the relation,
$\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}$ (176)
$=$ $\frac{1}{\not\!\!D_{\rm SO(3)}+m}+\frac{1}{\not\!\!D_{\rm SO(3)}+m}(-\mu\gamma_{4})\frac{1}{\not\!\!D_{\rm SO(3)}+m+\mu\gamma_{4}}$ (177)
$=$ $\frac{1}{\not\!\!D_{\rm SO(3)}+m}\sum_{k=0}^{\infty}\left(-\mu\gamma_{4}\frac{1}{\not\!\!D_{\rm SO(3)}+m}\right)^{k},$ (178)
further expansion of the series is possible for both $p(\mu)$ and $q(\mu)$,
$p(\mu)\equiv\sum_{n=1}^{\infty}\frac{(-)^{n+1}g^{2n}}{2n}{\rm Tr}\left[\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m}\sum_{k=0}^{\infty}\left(-\mu\gamma_{4}\frac{1}{\not\!\!D_{\rm SO(3)}+m}\right)^{k}\right]^{2n}\in{\bf R},$ (179)
$q(\mu)\equiv\sum_{n=0}^{\infty}\frac{(-)^{n+1}g^{2n+1}}{2n+1}{\rm Tr}\left[\not\!\!A_{\rm SU(3)/SO(3)}\frac{1}{\not\!\!D_{\rm SO(3)}+m}\sum_{k=0}^{\infty}\left(\mu\gamma_{4}\frac{1}{\not\!\!D_{\rm SO(3)}+m}\right)^{2k+1}\right]^{2n+1}\in{\bf R}.$ (180)
The series for even $k$ vanishes in the second equation due to a symmetric relation similar to Eq. (173). The second equation again shows that $q(\mu)$ is an odd function of $\mu$. This expansion is useful when $\mu$ is small, yet preserves the effect of $\mu$ in the SO(3) fermion determinant.
If the norm of $A^{\mu}_{\rm SU(3)/SO(3)}$ is well suppressed in the maximally SO(3) gauge, $q(\mu)$ may not fluctuate significantly over the sampled gauge configurations, and also the infinite series in Eq. (LABEL:eq:expansion) could be truncated to a finite one. Even after the truncation, the effect of the chemical potential $\mu$ is firmly included in the SO(3) fermionic determinant.
@@ VIII SO(3) real algebra method
Now, we formulate the SO(3) real algebra method in SU(3) lattice QCD at a finite density with a quark-number chemical potential $\mu$. Using the SU(3) lattice fermion kernel $K[U;\mu]$, the SU(3) fermionic determinant is expressed as ${\rm Det}~K[U;\mu]$. The SU(3) lattice QCD generating functional in the maximal SO(3) gauge at finite densities is expressed as
$Z^{\rm~SU(3)}_{\rm SO(3)~gauge}(\mu)$ $=$ $\int DUe^{-S_{G}[U]}~\delta^{\rm GF}_{\rm SO(3)}[U]~\Delta_{\rm SO(3)}^{\rm FP}[U]~{\rm Det}~K[U;\mu]$ (181)
@@
The next important step is to check the phase factor of the determinant ratio ${\rm Det}~K[U_{k};\mu]/{\rm Det}~K[u_{k};\mu]$ in the maximal SO(3) gauge in an actual lattice QCD Monte Carlo simulation. The degree of fluctuation would depend on the lattice volume, the chemical potential $\mu$, the fermion mass $m$ and so on. If there is a window that is close to the physical situation, this approach would be realistic for exploring finite density SU(3) QCD.
@@ IX Summary and conclusion
In this paper, we have proposed the SO(3) real algebra method for SU(3) lattice QCD calculations at finite baryon-number densities. In this method, we have divided the SU(3) gauge variable into the SO(3) and SU(3)/SO(3) parts. Then, we have introduced the maximal SO(3) gauge by minimizing the SU(3)/SO(3) part of the SU(3) gauge variable. We have considered the SO(3) fermionic determinant, that is, the fermionic determinant of the SO(3) part of the SU(3) gauge variable.
@@
The SO(3) real algebra method alternates between the maximal SO(3) gauge fixing and Monte Carlo updates on the SO(3) determinant and $e^{-S_{G}}$. After the most importance sampling, we have treated the remaining ratio of the SU(3) and SO(3) fermionic determinants as a weight factor in the reweighting method. In fact, we have only used gauge fixing for the most importance sampling, and the final result is fully gauge invariant. If the phase factor of the ratio does not fluctuate significantly in the sampled gauge configurations, then SU(3) lattice QCD calculations at finite densities would be feasible. This method can be extended to finite temperature QCD by imposing temporal (anti)periodicity. In fact, this method can be used to study QCD at finite densities and temperatures.
@@
If there is a window that is close to the physical situation, this approach would be realistic for lattice calculations of SU(3) QCD at finite densities and temperatures, and the QCD phase diagram would be obtained.
@@
Compared to other methods, the SO(3) real algebra method is a type of reweighting method [28, 29, 30]. This method is philosophically similar to the Lefschetz thimble method [36, 37, 38, 39, 40, 41], which reduces phase fluctuations by taking a suitable complex path, while this method reduces them by taking a suitable gauge. Unlike the Taylor expansion method [20, 21], this method does not use an expansion by $\mu$, so its applicable region is not limited to small $\mu$. Even if this method alone may not sufficiently suppress phase fluctuations, combining it with the other methods mentioned in Sec. I may enable the numerical simulation of finite density SU(3) QCD feasible.